a reaction takes place in 3 steps with rate constant k1, k2, k3. the overall rate rate constant is k= k1*(k2)^1/2 /k3. if the activation energies are 40, 30 ,20 kj for step 1,2,3 respectively then find the overall activation energy of the reaction. please tell me the solution of this answer as soon as possible.

 
 

Grade:

3 Answers

Naveen Kumar
9 years ago
Given k=k1*(k2)1/2/k3
Taking log both the side
lnk=lnk1+lnk2(1/2)-lnk3
We know that k=Ae-Ea/RT;lnk=lnA-Ea/RT
let every step have same arrhenius constant(i.e.A) and Ea1,Ea2 and Ea3 are the corresponding activation energy for 1st,2nd and the third step respectively.
So, k=lnA-Ea/RT=lnA-Ea1/RT..+(1/2)*[lnA-Ea2/RT]-lnA+Ea3/RT
So if we have value of A then we can get the value of Ea using the equation.
SIMRAN TAHLANI
7 years ago
Net Activation Energy= 40+ 1/2*30- 20=35 KJ. After Solving it and removing logA in the equation. Ea1+1/2*Ea2-Ea3
Ayush
6 years ago
,2nd and the third step respectively.So, k=lnA-Ea/RT=lnA-Ea1/RT..+(1/2)*[lnA-Ea2/RT]-lnA+Ea3/RTSo if we have value of A then we can get the value of Ea using the equationst;lnk=lnA-Ea/RTlet every step have same arrhenius constant(i.e.A) and Ea1,Ea2 and Ea3 are the corresponding activation energy for 1-Ea/RT/k3Taking log both the sidelnk=lnk1+lnk2(1/2)-lnk3We know that k=Ae1/2Given k=k1*(k2)